# A Step to Decouple Optimization in 3DGS

### v1

@@ 5 Experiments
Dataset and Metric Following existing research (kerbl20233d, ), we employ 13 scenes from 3 datasets, including 9 scenes in Mip-Nerf360 (barron2022mip, ), 2 scenes from the Tanks&Temples dataset (knapitsch2017tanks, ) and 2 scenes provided by Hedman et al. (hedman2018deep, ).
We use a train/test split for datasets, using the methodology suggested by Mip-NeRF360, taking every 8th photo for test, for consistent and meaningful comparisons to generate the error metrics, using the standard PSNR, SSIM, and L-PIPS(vgg) metrics used most frequently in the literature. We also present the total primitive number , the active primitives number and the dead primitives number . Given the significant variation in primitive number across methods (often differing by millions), we employ normalized changes to the vanilla (, where ) to better illustrate primitive number variations. The active/dead primitive is defined by opacity larger than 1/255 or not. This definition follows the underlying 3DGS333https://github.com/graphdeco-inria/gaussian-splatting implementation, where primitives with opacity below 1/255 are excluded from rendering, differing from (kheradmand20243d, ).
Baselines To evaluate the impact of different components on the training pipeline, we conduct extensive experiments on both the vanilla 3DGS and 3DGS-MCMC frameworks. These two baselines respectively represent approaches without or with an maximum. Without explicitly introducing pruning operations, we observe that a substantial number of redundant primitives can be automatically removed under the vanilla 3DGS with AdamW-GS. For comparison, we include two adaptive pruning methods, learning-based MaskGaussion (liu2024maskgaussian, ) and Redundant Primitive Removal (RePR) (papantonakis2024reducing, ) with hand-crafted criterion. Unlike approaches with a predefined pruning rate, these methods automatically identify redundant primitives and claim to achieve this with negligible performance degradation. All methods are run on a single A6000 GPU.
@@ 5.1 Results and Analysis
Activated via the Moment Within the 3DGS-MCMC framework, we evaluate the effect of our proposed RSR on regularization, a way of activating regularization via rescaling the moment. Regardless of the original L1 form subject to gradient coupling or our proposed DAR, attribute regularization remains tied to the second moment. Figure 3 e–f shows that applying RSR substantially increases the number of dead primitives, thereby strengthening attribute regularization. This directly demonstrates the effectiveness of RSR as a component to amplify regularization. However, as highlighted by the comparison between MCMC-Sparse-RSR and MCMC-AdamWGS-MC17 in Figure 3 f, the influence of RSR with same StSS is weaker under L1 regularization with gradient coupling than under DAR. By decoupling the gradient in the moment and recomposing the attribute regularization, RSR more effectively amplifies its impact, which also translates into improved reconstruction quality under stronger regularization.
Side Effect in Coupling To further examine this effect, we increase the sampling ratio of RSR in MidNeRF-360 indoor scenes (which typically involve fewer primitives), with results summarized in Appendix, from Sec. I.6 to Sec. I.9, MC20 and MC21, in which we replace the low-ratio StSSMC1 in original 3DGSMCMC with L1 regularization and 3DGSMCMC with our proposed AdamW-GS with higher-ratio StSSMC2. As the sampling ratio increases, original 3DGS-MCMC with L1 regularization exhibits significant drops in reconstruction quality across all four scenes (e.g., Room: 32.514→31.179 dB; Kitchen: 32.289→31.924 dB). The same phenomenon is also observed in the experiments on outdoor scenes. In contrast, 3DGS-MCMC with DAR shows little to no degradation in Room, Counter, and Bosai, and only a minor drop in Kitchen (32.546→32.298 dB), while consistently outperforming L1 regularization under the same settings. These results confirm that coupling hinders effective control of regularization and can easily introduce negative effects on reconstruction quality. Thus, decoupling the gradient from the moment and recomposing regularization is necessary for stable and effective optimization.
@@
We compare vanilla 3DGS trained with AdamW-GS against two adaptive pruning methods. Importantly, our approach contains no extra pruning component; instead, it relies solely on opacity DAR. In terms of pruning performance, AdamW-GS achieves results comparable to MaskGaussion and even outperforms it by 2% on outdoor scenes. Beyond pruning efficiency, our method also improves reconstruction quality: in outdoor scenes, it reduces the number of primitives by 48.4% while increasing PSNR by 0.2 dB and SSIM by 0.01. In indoor scenes, it removes 50% of primitives while still improving PSNR by 0.1 dB. By contrast, MaskGaussion removes 61% of primitives but degrades PSNR by 0.1 dB, which is because MaskGaussion suffers from potential reconstruction quality risk (see Figure 2 c). Visualization in Figure 3 c-d also demonstrates better detail reconstruction in our methods compared to vanilla 3DGS, which is one of the reasons for relatively more primitives than MaskGaussion. Ultimately, our objective is not simply to minimize the number of active primitives, but to strike a preferable balance—achieving higher reconstruction quality with as few primitives as necessary.
@@
| MaskGaussion | 29.895 | 0.908 | 0.248 | -65% | 23.607 | 0.846 | 0.181 | -50% |
Additional datasets are also evaluated with our proposed method. As shown in Table 4, experiments on Deep Blending and Tanks&Temples further demonstrate the superiority of our approach. On Deep Blending in particular, 3DGS equipped with AdamW-GS even surpasses the PSNR of the original 3DGS-MCMC. In Appendix Sec. C.4, we also report results on long-sequence datasets for vanilla 3DGS and 3DGS-MCMC, both with and without our proposed AdamW-GS, where our method continues to show clear advantages.
@@
To further validate our findings and the robustness of our method, we apply AdamW-GS to additional pipelines. As shown in Table 3 and discussed therein, MaskGaussian exhibits a potential reconstruction-quality risk on indoor datasets—a phenomenon consistent with Observation 1 in Sec. 4.1 from the Adam vs. Sparse Adam experiments. We attribute this issue to the synchronous updating of mask scores in MaskGaussian (a brief introduction to MaskGaussion is given in Sec.C). Results in Table 5 further support this hypothesis: when trained with AdamW-GS, MaskGaussian no longer suffers from the quality risk and additionally prunes about 7% redundant primitives on indoor scenes. This also demonstrates that our method is compatible with additional pruning operations. In Appendix Sec. C, we provide results from more pipeline variants combined with AdamW-GS, offering further evidence of the robustness of our approach.
@@
To better understand how our method penalizes redundant primitives, we further visualize the primitive number dynamics during training, as shown in Figure 2. PePR, which relies solely on opacity L1 regularization during densification, shows limit in reducing redundancy. In contrast, with our proposed regularization, the number of primitives remains consistently lower. Even in cases where is larger than that of MaskGaussian, our method (vanilla 3DGS trained with AdamW-GS)—restricted to the densification stage—still outperforms MaskGaussian in reducing redundancy during densification.
@@
Extra Exploration is Necessary Previous studies have highlighted the role of exploration in 3DGS(jung2024relaxing, ; kheradmand20243d, ), and here we investigate its interplay with redundancy. AdamW-GS benefits from the noise-based position regularization, yielding improved pruning efficiency and higher SSIM. As shown in Table 2 GS7 and GS8, removing opacity reset weakens the noise-based regularization, leading to less improvement in SSIM and redundancy removal. We attribute the effectiveness of the noise-based regularization in AdamW-GS to its more appropriate attribute regularization. Overall, our results underscore that additional exploration is necessary. However, the noise-based position regularization from 3DGS-MCMC cannot be directly applied to scenes with relatively few primitives (e.g., indoor scenes), where their higher sensitivity to noise results in a drop in reconstruction quality. In Appendix Sec. D, we provide a more detailed analysis of the exploration ability induced by our method. We also introduce two additional exploration strategies that can be combined with current AdamW-GS and yield further improvements on certain indoor scenes as well.
@@
Efficiency Analysis Benefiting from Sparse Adam, as shown in Table 3, 3DGS-MCMC further improves optimization efficiency. Vanilla 3DGS enjoys the same benefit, while our method achieves additional gains by substantially reducing redundancy during the densification stage, thereby further enhancing optimization efficiency. More detailed analysis are provided in Appendix Sec.A.
@@ 6 Discussion and Conclusion
This work take a step to decouple the complex 3DGS optimization and analyzes key mechanisms. AdamW-GS is further proposed, which improves efficiency and enables more controllable regularization. The method enhances reconstruction quality in the MCMC framework and achieves notable redundancy removal in vanilla 3DGS with slight quality gains and even without pruning. Our results indicate that efficiency and effectiveness can be jointly improved given appropriate regularization. While we focused on opacity and scaling, more common DAR in different attribute or alternative strategies may yield further improvements. Despite the robust current hyperparameters and hand-crafted SStS, changing the parameters during training in an adaptive manner may offer a more promising direction. Finally, although noise-driven exploration shows necessary, it can harm certain scenes. Other provided exploration strategies offer limited improvement. A more systematic investigation of exploration strategies may be a promising direction for future research.
Appendix A Time Cost Analysis
@@ Appendix B Preliminary
3DGS approximates the radiance field of targeted scenes via a group of Gaussion primitives parameterized with where , denotes the primitive position, is the positive semi-definite covariance matrix decoupled into quaternion and a scaling vector, is considered as the opacity value. The color attribute is typically stored as spherical harmonics coefficients and converted into RGB when rendering.
Given a viewpoint and transformation [1], the pixel on the screen is rendered via -blending according to the projected depth of primitives:
@@
The 3DGS training pipeline typically initializes from Structure-from-Motion (SfM) [33] followed by several warm-up iterations [1, 29]. The primitives then undergo densification, wherein new primitives are generated to enhance scene representation while redundant ones are pruned. In the vanilla pipeline (Figure 9 a), both operations are governed by adaptive density control: new primitives are created via cloning or splitting based on gradient magnitude, whereas low-opacity primitives are removed. After densification, optimization proceeds in a pure optimization (P-Op) stage, during which only gradient propagation and attribute updates occur. Throughout warm-up, densification, and P-Op, training employs a photometric loss (left in Eq.1) with the Adam optimizer. Building on Stochastic Gradient Langevin Dynamics (SGLD), 3DGS-MCMC [10] follows a similar stage division, as shown in Figure 9 b: new primitive sampling and dead primitive reallocation during densification, followed by a shortened P-Op. It also adopts Adam, augmented with a regularization loss in Eq. 1 to promote respawn and additional positional noise for exploration or adherence to the SGLD form.
Appendix C More Pipeline Varients with AdamW-GS or Experiments on Different Datasets
@@
C.1 MaskGaussion with AdamW-GS: More Stable Updating of Mask Score
MaskGaussian introduces a probabilistic formulation of 3D Gaussian primitives, where each primitive is assigned a learnable probability of existence that governs a dynamic sampling process during rendering. This mechanism enables adaptive pruning of redundant primitives throughout the 3DGS optimization. Specifically, each primitive is associated with a mask score , from which a binary mask is stochastically sampled using the Gumbel-Softmax reparameterization. The standard rendering equation in Eq. 9 is thus modified to incorporate the mask in both color accumulation and transmittance updates, as shown in Eq. 10. Detailed derivations and implementation are provided in [12].
@@ C.1 MaskGaussion with AdamW-GS: More Stable Updating of Mask Score
C.2 Taming-3DGS with AdamW-GS
@@
Taming-3DGS [7] constrains the total number of primitives by restricting the number of new primitives added at each densification step, as defined by Eq. 12. During every densification step, Taming-3DGS computes a score for each primitive to determine its probability of being selected for densification. Based on these scores, a fixed number of primitives are sampled for densification. Additionally, Taming-3DGS incorporates further optimizations, such as enhanced parallelization. For detailed implementation and algorithmic design, please refer to [7].
@@
Two operating modes are provided in Taming-3DGS444https://github.com/humansensinglab/taming-3dgs: final count and multiplier, which correspond to a higher and a lower upper bound on the final total primitive umber , respectively. Experiments were conducted under both modes for Taming-3DGS as well as Taming-3DGS with AdamW-GS. DAR in this setting only includes opacity and scaling regularization as well, following the same configuration as in Sec. 4.4. Noise regularization is omitted because the growth rate in Taming-3DGS is constrained. Considering that AdamW-GS imposes stronger penalties on redundant primitives, leading to the generation of a large number of dead primitives that are rapidly pruned and thus reduce the total primitive count significantly, we propose to replace the conservative pruning strategy in Taming-3DGS with the pruning method used in vanilla 3DGS when integrating AdamW-GS. We denote this modified pipeline as Taming-3DGS-p. Related experimental results are summarized in Table 8.
@@
C.3 Deformable Beta Splatting with AdamW-GS

### v3

@@ 5 Experiments
Dataset and Metric Following existing research (Kerbl et al., 2023), we employ 13 scenes from 3 datasets, including 9 scenes in Mip-Nerf360 (Barron et al., 2022), 2 scenes from the Tanks&Temples dataset (Knapitsch et al., 2017) and 2 scenes provided by Hedman et al. (Hedman et al., 2018).
We use a train/test split for datasets, using the methodology suggested by Mip-NeRF360, taking every 8th photo for test, for consistent and meaningful comparisons to generate the error metrics, using the standard PSNR, SSIM, and L-PIPS(vgg) metrics used most frequently in the literature. We also present the total primitive number , the active primitives number and the dead primitives number . Given the significant variation in primitive number across methods (often differing by millions), we employ normalized changes to the vanilla (, where ) to better illustrate primitive number variations. The active/dead primitive is defined by opacity larger than 1/255 or not. This definition follows the underlying 3DGS333https://github.com/graphdeco-inria/gaussian-splatting implementation, where primitives with opacity below 1/255 are excluded from rendering, differing from (Kheradmand et al., 2024).
Baselines To evaluate the impact of different components on the training pipeline, we conduct extensive experiments on both the vanilla 3DGS and 3DGS-MCMC frameworks. These two baselines respectively represent approaches without or with an maximum. Without explicitly introducing pruning operations, we observe that a substantial number of redundant primitives can be automatically removed under the vanilla 3DGS with AdamW-GS. For comparison, we include two adaptive pruning methods, learning-based MaskGaussian (Liu et al., 2025b) and Redundant Primitive Removal (RePR) (Papantonakis et al., 2024) with hand-crafted criterion. Unlike approaches with a predefined pruning rate, these methods automatically identify redundant primitives and claim to achieve this with negligible performance degradation. All methods are run on a single A6000 GPU.
@@ 5.1 Results and Analysis
Activated via the Moment Within the 3DGS-MCMC framework, we evaluate the effect of our proposed RSR on regularization, a way of activating regularization via rescaling the moment. Regardless of the original L1 form subject to gradient coupling or our proposed DAR, attribute regularization remains tied to the second moment. Figure 3 e–f shows that applying RSR substantially increases the number of dead primitives, thereby strengthening attribute regularization. This directly demonstrates the effectiveness of RSR as a component to amplify regularization. However, as highlighted by the comparison between MCMC-Sparse-RSR and MCMC-AdamWGS-MC17 in Figure 3 e, the influence of RSR with same StSS is weaker under L1 regularization with gradient coupling than under DAR. By decoupling the gradient in the moment and recomposing the attribute regularization, RSR more effectively amplifies its impact, which also translates into improved reconstruction quality under stronger regularization.
Side Effect in Coupling To further examine this effect, we increase the sampling ratio of RSR in MidNeRF-360 indoor scenes (which typically involve fewer primitives), with results summarized in Appendix, from Sec. K.6 to Sec. K.9, MC20 and MC21, in which we replace the low-ratio StSSMC1 in original 3DGSMCMC with L1 regularization and 3DGSMCMC with our proposed AdamW-GS with higher-ratio StSSMC2. As the sampling ratio increases, original 3DGS-MCMC with L1 regularization exhibits significant drops in reconstruction quality across all four scenes (e.g., Room: 32.514→31.179 dB; Kitchen: 32.289→31.924 dB). The same phenomenon is also observed in the experiments on outdoor scenes. In contrast, 3DGS-MCMC with DAR shows little to no degradation in Room, Counter, and Bosai, and only a minor drop in Kitchen (32.546→32.298 dB), while consistently outperforming L1 regularization under the same settings. These results confirm that coupling hinders effective control of regularization and can easily introduce negative effects on reconstruction quality. Thus, decoupling the gradient from the moment and recomposing regularization is necessary for stable and effective optimization.
@@
We compare vanilla 3DGS trained with AdamW-GS against two adaptive pruning methods. Importantly, our approach contains no extra pruning component; instead, it relies solely on opacity DAR. In terms of pruning performance, AdamW-GS achieves results comparable to MaskGaussian and even outperforms it by 2% on outdoor scenes. Beyond pruning efficiency, our method also improves reconstruction quality: in outdoor scenes, it reduces the number of primitives by 48.4% while increasing PSNR by 0.2 dB and SSIM by 0.01. In indoor scenes, it removes 50% of primitives while still improving PSNR by 0.1 dB. By contrast, MaskGaussian removes 61% of primitives but degrades PSNR by 0.1 dB, which is because MaskGaussian suffers from potential reconstruction quality risk (see Figure 2 c). Visualization in Figure 3 c-d also demonstrates better detail reconstruction in our methods compared to vanilla 3DGS, which is one of the reasons for relatively more primitives than MaskGaussian. Ultimately, our objective is not simply to minimize the number of active primitives, but to strike a preferable balance—achieving higher reconstruction quality with as few primitives as necessary.
@@
| MaskGaussian | 29.895 | 0.908 | 0.248 | -65% | 23.607 | 0.846 | 0.181 | -50% |
Additional datasets are also evaluated with our proposed method. As shown in Table 4, experiments on Deep Blending and Tanks&Temples further demonstrate the superiority of our approach. On Deep Blending in particular, 3DGS equipped with AdamW-GS even surpasses the PSNR of the original 3DGS-MCMC. In Appendix Sec. E.4, we also report results on long-sequence datasets for vanilla 3DGS and 3DGS-MCMC, both with and without our proposed AdamW-GS, where our method continues to show clear advantages.
@@
To further validate our findings and the robustness of our method, we apply AdamW-GS to additional pipelines. As shown in Table 3 and discussed therein, MaskGaussian exhibits a potential reconstruction-quality risk on indoor datasets—a phenomenon consistent with Observation 1 in Sec. 4.1 from the Adam vs. Sparse Adam experiments. We attribute this issue to the synchronous updating of mask scores in MaskGaussian (a brief introduction to MaskGaussian is given in Sec. E). Results in Table 5 further support this hypothesis: when trained with AdamW-GS, MaskGaussian no longer suffers from the quality risk and additionally prunes about 7% redundant primitives on indoor scenes. This also demonstrates that our method is compatible with additional pruning operations. In Appendix Sec. E, we provide results of more pipeline variants with AdamW-GS, offering further evidence of the robustness.
@@
To better understand how our method penalizes redundant primitives, we further visualize the primitive number dynamics during training, as shown in Figure 2. PePR, which relies solely on opacity L1 regularization during densification, shows limit in reducing redundancy. In contrast, with our proposed regularization, the number of primitives remains consistently lower. Even in cases where is larger than that of MaskGaussian, our method (vanilla 3DGS with AdamW-GS)—restricted to the densification—still outperforms MaskGaussian in reducing redundancy during densification.
@@
Extra Exploration is Necessary Previous studies have highlighted the role of exploration in 3DGS (Jung et al., 2024; Kheradmand et al., 2024), and here we investigate its interplay with redundancy. AdamW-GS benefits from the noise-based position regularization, yielding improved pruning efficiency and higher SSIM. As shown in Table 2 GS7 and GS8, removing opacity reset weakens the noise-based regularization, leading to less improvement in SSIM and redundancy removal. We attribute the effectiveness of the noise-based regularization in AdamW-GS to its more appropriate attribute regularization. Overall, our results underscore that additional exploration is necessary. However, the noise-based position regularization from 3DGS-MCMC cannot be directly applied to scenes with relatively few primitives (e.g., indoor scenes), where their higher sensitivity to noise results in a drop in reconstruction quality. In Appendix Sec. F, we provide a more detailed analysis of the exploration ability induced by our method. We also introduce two additional exploration strategies that can be combined with current AdamW-GS and yield further improvements on certain indoor scenes as well.
@@
Efficiency Analysis Benefiting from Sparse Adam, as shown in Table 3, 3DGS-MCMC further improves optimization efficiency. Vanilla 3DGS enjoys the same benefit, while our method achieves additional gains by substantially reducing redundancy during the densification stage, thereby further enhancing optimization efficiency. More detailed analysis is provided in Appendix Sec. C.
@@ 6 Discussion and Conclusion
This work takes a step to decouple the complex 3DGS optimization and analyzes key mechanisms. AdamW-GS is further proposed, which improves efficiency and enables more controllable regularization. The method enhances reconstruction quality in the MCMC framework and achieves notable redundancy removal in vanilla 3DGS with slight quality gains and even without pruning. Our results indicate that efficiency and effectiveness can be jointly improved given appropriate regularization. While we focused on opacity and scaling, more common DAR in different attribute or alternative strategies may yield further improvements. Despite the robust current hyperparameters and hand-crafted SStS, changing the parameters during training in an adaptive manner may offer a more promising direction. Finally, although noise-driven exploration shows necessary, it can harm certain scenes. Other provided exploration strategies offer limited improvement. A more systematic investigation of exploration strategies may be a promising direction for future research.
Acknowledgments This work was supported by China Mobile Hunan Company Limited, China Mobile Communications Group Co., Ltd., and by the National Natural Science Foundation of China under Grant U22B2050, 62425305, 62221002, and 62503161. This work was conducted as part of the project “Networked Robotic System for Major Equipment Manufacturing (5G+Robotics)”.
Appendix A Reproducibility
The datasets are directly available from (Kerbl et al., 2023). Methodological steps and formulas are detailed in Secs. 4.4 and J.1, while all hyperparameters are reported in Secs. J.2 and K.
More rendering visualization can be found in Figure 10.
Appendix B The Use of Large Language Models (LLMs)
In this work, LLMs are used solely to assist in writing and polishing the manuscript; they do not contribute to the discovery or experimental design.
Appendix C Time Cost Analysis
@@ Appendix D Preliminary
3DGS approximates the radiance field of targeted scenes via a group of Gaussian primitives parameterized with where , denotes the primitive position, is the positive semi-definite covariance matrix decoupled into quaternion and a scaling vector, is considered as the opacity value. The color attribute is typically stored as spherical harmonics coefficients and converted into RGB when rendering.
Given a viewpoint and transformation (Kerbl et al., 2023), the pixel on the screen is rendered via -blending according to the projected depth of primitives:
@@
The 3DGS training pipeline typically initializes from Structure-from-Motion (SfM) (Schonberger & Frahm, 2016) followed by several warm-up iterations (Kerbl et al., 2023; Jung et al., 2024). The primitives then undergo densification, wherein new primitives are generated to enhance scene representation while redundant ones are pruned. In the vanilla pipeline (Figure 9 a), both operations are governed by adaptive density control: new primitives are created via cloning or splitting based on gradient magnitude, whereas low-opacity primitives are removed. After densification, optimization proceeds in a pure optimization (P-Op) stage, during which only gradient propagation and attribute updates occur. Throughout warm-up, densification, and P-Op, training employs a photometric loss (left in Eq. 1) with the Adam optimizer. Building on Stochastic Gradient Langevin Dynamics (SGLD), 3DGS-MCMC (Kheradmand et al., 2024) follows a similar stage division, as shown in Figure 9 b: new primitive sampling and dead primitive reallocation during densification, followed by a shortened P-Op. It also adopts Adam, augmented with a regularization loss in Eq. 1 to promote respawn and additional positional noise for exploration or adherence to the SGLD form.
Appendix E More Pipeline Variants with AdamW-GS or Experiments on Different Datasets
@@
E.1 MaskGaussian with AdamW-GS: More Stable Updating of Mask Score
MaskGaussian introduces a probabilistic formulation of 3D Gaussian primitives, where each primitive is assigned a learnable probability of existence that governs a dynamic sampling process during rendering. This mechanism enables adaptive pruning of redundant primitives throughout the 3DGS optimization. Specifically, each primitive is associated with a mask score , from which a binary mask is stochastically sampled using the Gumbel-Softmax reparameterization. The standard rendering equation in Eq. 9 is thus modified to incorporate the mask in both color accumulation and transmittance updates, as shown in Eq. 10. Detailed derivations and implementation are provided in (Liu et al., 2025b).
@@ E.1 MaskGaussian with AdamW-GS: More Stable Updating of Mask Score
E.2 Taming-3DGS with AdamW-GS
@@
Taming-3DGS (Mallick et al., 2024) constrains the total number of primitives by restricting the number of new primitives added at each densification step, as defined by Eq. 12. During every densification step, Taming-3DGS computes a score for each primitive to determine its probability of being selected for densification. Based on these scores, a fixed number of primitives are sampled for densification. Additionally, Taming-3DGS incorporates further optimizations, such as enhanced parallelization. For detailed implementation and algorithmic design, please refer to (Mallick et al., 2024).
@@
Two operating modes are provided in Taming-3DGS444https://github.com/humansensinglab/taming-3dgs: final count and multiplier, which correspond to a higher and a lower upper bound on the final total primitive number , respectively. Experiments were conducted under both modes for Taming-3DGS as well as Taming-3DGS with AdamW-GS. DAR in this setting only includes opacity and scaling regularization as well, following the same configuration as in Sec. 4.4. Noise regularization is omitted because the growth rate in Taming-3DGS is constrained. Considering that AdamW-GS imposes stronger penalties on redundant primitives, leading to the generation of a large number of dead primitives that are rapidly pruned and thus reduce the total primitive count significantly, we propose to replace the conservative pruning strategy in Taming-3DGS with the pruning method used in vanilla 3DGS when integrating AdamW-GS. We denote this modified pipeline as Taming-3DGS-p. Related experimental results are summarized in Table 8.
@@
E.3 Deformable Beta Splatting with AdamW-GS